# Quantum field theory treatment of oscillations of Dirac neutrinos in external fields

### v1

@@ 5.4 QFT description of the spin-flavor precession in a magnetic field
We have derived all kinds of transition probabilities for spin-flavor precession of Dirac neutrinos in Sec. 5.2. Despite the transitions like $\nu_{e\mathrm{L}}\to\nu_{(e,\mu)\mathrm{R}}$ are formally allowed, see Eqs. (5.28) and (5.29), we cannot detect right-handed Dirac neutrinos. We assume that a neutrino detector is based on the standard model physics, i.e. it involves the leptonic current in Eq. (3.2). This current contains only the left-handed neutrino fields. That is why the matrix elements for the processes $\nu_{e\mathrm{L}}\to\nu_{\mu\mathrm{R}}$ and $\nu_{e\mathrm{L}}\to\nu_{e\mathrm{R}}$ are zero in our model if we study ultrarelativistic neutrinos.
It is the key difference between the spin-flavor precession of Majorana and Dirac neutrinos. We recall that the marker of the spin-flavor precession in the Majorana case is the appearance of an anti-charged-lepton. That is, if we consider the process $\nu_{e}\xrightarrow{B}\bar{\nu}_{\mu}$ for Majorana $\nu_{e}$ and $\nu_{\mu}$, one has an electron $e^{-}$ in a source and antimuon $\mu^{+}$ in a detector [18]. Therefore, to study the impact of the magnetic field on the spin-flavor precession of Dirac neutrinos in frames of the QFT based approach, one has to consider the transitions $\nu_{e\mathrm{L}}\to\nu_{(e,\mu)\mathrm{L}}$ and try to rederive Eqs. (5.26) and (5.27). Since we deal with only left particles in the initial and final states, the Feynman diagrams in Fig. 2 in the magnetic and matter cases formally coincide.
@@
In our calculations, we adopt the same geometry of the system as in Sec. 4.3. Namely, $\mathbf{L}=L\mathbf{e}_{z}$. Analogously to Sec. 5.2, we take that the magnetic field is transverse to the line connecting the source and the detector, $\mathbf{B}=B\mathbf{e}_{x}$. Then, we also assume the forward scattering approximation for charged leptons, i.e., $u_{e,\mu}^{(\mathrm{L})\mathrm{T}}=(0,\kappa_{e,\mu})$, where $\kappa_{e,\mu}^{\mathrm{T}}=(0,1)$.
@@
$\mathcal{M}_{e\to\mu}=-\frac{\mathrm{i}Ee^{\mathrm{i}EL}}{2\pi L}\frac{\sin 2\theta\frac{\Delta m^{2}}{4E}\sin\mathfrak{E}_{\mathrm{B}}L}{\mathfrak{E}_{\mathrm{B}}}\left[1-\frac{(\mu B)^{2}}{2E\mathfrak{E}_{\mathrm{B}}}\right],$ (5.35)
@@
Analogously to Eq. (4.27), Eq. (5.35) is valid for for ultrarelativistic neutrinos and applicable for sufficiently long propagation distances $L\gg E^{-1}$.
@@
where $P_{\mathrm{max}}\approx P_{\nu_{e\mathrm{L}}\to\nu_{\mu\mathrm{L}}}^{(\mathrm{qm})}+P_{\nu_{e\mathrm{L}}\to\nu_{\mu\mathrm{L}}}^{(\mathrm{qft})}$, and
@@
$P_{\nu_{e\mathrm{L}}\to\nu_{\mu\mathrm{L}}}^{(\mathrm{qft})}$ $=-\sin^{2}2\theta\frac{\left(\frac{\Delta m^{2}}{4E}\right)^{2}(\mu B)^{2}}{E\left[(\mu B)^{2}+\left(\frac{\Delta m^{2}}{4E}\right)^{2}\right]^{3/2}}.$ (5.39)
@@
The correction to the transition probability $P_{\nu_{e\mathrm{L}}\to\nu_{\mu\mathrm{L}}}^{(\mathrm{qft})}$ in Eq. (5.39) arises from the term $\propto\mu^{2}(\mathbf{B}\cdot\hat{p})^{2}$ in the propagators in Eqs. (5.32) and (5.33), which is discussed at the end of Sec. 5.3. One can see that $P_{\nu_{e\mathrm{L}}\to\nu_{\mu\mathrm{L}}}^{(\mathrm{qft})}<0$, i.e. the total transition probability is less than the QM prediction. Moreover, one gets from Eqs. (5.38) and (5.39) that $|P_{\nu_{e\mathrm{L}}\to\nu_{\mu\mathrm{L}}}^{(\mathrm{qft})}|/P_{\nu_{e\mathrm{L}}\to\nu_{\mu\mathrm{L}}}^{(\mathrm{qm})}\sim(\mu B)^{2}/\mathfrak{E}_{\mathrm{B}}E\sim\mu B/E\ll 1$ provided that the magnetic contribution is dominant. Therefore, the QFT correction to the propagators, $\propto\mu^{2}(\mathbf{B}\cdot\hat{p})^{2}$, can be safely neglected since its contribution to the transition probability is negligible. We also mention that the QFT correction to the probability of the spin-flavor precession of Majorana neutrinos, $\nu_{e}\xrightarrow{B}\bar{\nu}_{\mu}$, analogous to that in Eq. (5.39), was obtained in Ref. [18].
@@
One can see in Eq. (5.40) that the leading term reproduces the QM result in Eq. (5.26). A small quantum correction $\sim\mu B/E$ to the amplitude of the survival probability is also present Eq. (5.40).
6 Conclusion
@@
The technique for finding of the exact propagators is solving the system of the Dyson equations, which are equivalent to summing of infinite series of Feynman diagrams, depicted in Fig. 2. In these diagrams, the nondiagonal interaction with matter is accounted for. This technique was elaborated in Ref. [17] where oscillations of Majorana neutrinos in matter were studied. In case of Dirac neutrinos, the effective matter potential is a singular operator. Since the reciprocal of this operator is involved in the Dyson equations, the straightforward solution of these equations is ambiguous. Nevertheless, we have regularized this potential in Sec. 4.2. After removing the regularization, we have obtained the dressed propagators in background matter in Eqs. (4.17) and (4.20). These propagators have been used in the matrix element to get the transition probability of flavor oscillations in matter in Eq. (4.3) which coincides with the prediction of the QM approach.
@@
Then, we have studied the spin-flavor precession of Dirac neutrinos in a magnetic field in Sec. 5. We have adopted the situation when one has only a transition magnetic moment in the mass basis of two neutrinos. This kind of magnetic moments is disfavored in some models. Nevertheless, such an external magnetic field maximally mixes different neutrino mass states. Thus, the QFT formalism for neutrino oscillations based on dressed propagators is highlighted. Moreover, one can consider arbitrary magnetic moments from the phenomenological point of view.
@@
We have revealed a peculiarity of the treatment of the spin-flavor precession of Dirac neutrinos in frames of QFT. This formalism involves both a source, a neutrino propagation, and a detector as inherent parts. Since our model of a detector is based on the standard model physics, where right-handed neutrinos are sterile, we cannot describe the channel like $\nu_{e\mathrm{L}}\to\nu_{\mu\mathrm{R}}$. From the mathematical point of view, the matrix element for $\nu_{e\mathrm{L}}\to\nu_{\mu\mathrm{R}}$ is zero since the leptonic current in Eq. (3.2) is based on left chiral projections. On the contrary, the process like $\nu_{e\mathrm{L}}\to\nu_{\mu\mathrm{R}}$ is allowed in QM description which is provided in Sec. 5.2; cf. Eq. (5.29).
@@
It is the main difference between the Dirac and Majorana cases. In the latter situation, the process like $\nu_{e}\to\bar{\nu}_{\mu}$ is well possible in the QFT formalism. The marker of such a spin-flavor precession is the appearance of an antimuon in a detector.
@@
One can see in the propagators in Eqs. (5.32) and (5.33) that there is a quantum term $\propto\mu^{2}(\mathbf{B}\cdot\hat{p})^{2}$ which cannot be neglected if we are in frames of QFT. We have derived the matrix element and the transition probability based on these propagators in Sec. 5.4. The leading term in the transition probability in Eq. (5.38) reproduces the QM result in Eq. (5.27). The correction to the QM expression is given in Eq. (5.39). It arises from the quantum term $\propto\mu^{2}(\mathbf{B}\cdot\hat{p})^{2}$ in the propagators. This correction turns out to be small for ultrarelativistic neutrinos for all reasonable parameters. Thus, we have revealed a posteriori, i.e. after the calculation of a transition probability, that the virtuality of massive neutrinos in propagators is negligible. Note that the correction to the transition probability, analogous to that in Eq. (5.39), was derived in Ref. [18] while applying QFT for the description of the spin-flavor precesion of Majorana neutrinos.
@@
Finally, we note that the impact of external fields on neutrino oscillations is an essentially nonperturbative phenomenon which requires the summation of infinite number of Feynman diagrams Fig. 2. Only in this case, one gets the correct form of dressed propagators exactly accounting for an external field, which result in the transition probabilities consistent with the QM predictions.
At the end of this section, we list the main assumptions made in the course of applying of QFT for neutrino oscillations in external fields. The ultrarelativity of neutrinos, the consideration of two mass eigenstates, and the spatial homogeneity of external fields were important in our analysis.

### v2

@@ 5.4 QFT description of the spin-flavor precession in a magnetic field
We have derived all kinds of transition probabilities for spin-flavor precession of Dirac neutrinos in Sec. 5.2. Despite the transitions like $\nu_{e\mathrm{L}}\to\nu_{(e,\mu)\mathrm{R}}$ are formally allowed, see Eqs. (5.28) and (5.29), we cannot detect right-handed Dirac neutrinos. We assume that a neutrino detector is based on the standard model physics, i.e. it involves the leptonic current in Eq. (3.2). This current contains only the left-handed neutrino fields. That is why the processes $\nu_{e\mathrm{L}}\to\nu_{\mu\mathrm{R}}$ and $\nu_{e\mathrm{L}}\to\nu_{e\mathrm{R}}$ cannot be considered in frames of the adopted model for the description of oscillations if one deals with ultrarelativistic neutrinos.
It is the key difference between the spin-flavor precession of Majorana and Dirac neutrinos. We recall that the marker of the spin-flavor precession in the Majorana case is the appearance of an anti-charged-lepton. That is, if we consider the process $\nu_{e}\xrightarrow{B}\bar{\nu}_{\mu}$ for Majorana $\nu_{e}$ and $\nu_{\mu}$, one has an electron $e^{-}$ in a source and antimuon $\mu^{+}$ in a detector [18]. Therefore, to study the impact of the magnetic field on the spin-flavor precession of Dirac neutrinos in frames of the QFT based approach, one has to consider the transitions $\nu_{e\mathrm{L}}\to\nu_{(e,\mu)\mathrm{L}}$ and try to rederive Eqs. (5.26) and (5.27). Since, in the initial and final states, we deal with left particles only, the Feynman diagrams in Fig. 2 in the magnetic and matter cases formally coincide.
@@
In our calculations, we adopt the same geometry of the system as in Sec. 4.3. Namely, $\mathbf{L}=L\mathbf{e}_{z}$. Analogously to Sec. 5.2, we take that the magnetic field is transverse to the line connecting the source and the detector, $\mathbf{B}=B\mathbf{e}_{x}$. Then, we also assume the forward scattering approximation for charged leptons, i.e., $u_{e,\mu}^{\mathrm{T}}=(0,0,0,1)$.
@@
$\mathcal{M}_{e\to\mu}=-\frac{\mathrm{i}Ee^{\mathrm{i}EL}}{2\pi L}\frac{\sin 2\theta\frac{\Delta m^{2}}{4E}\sin\mathfrak{E}_{\mathrm{B}}L}{\mathfrak{E}_{\mathrm{B}}}\left(1-\frac{(\mu B)^{2}}{2E\mathfrak{E}_{\mathrm{B}}}\right)\left[1+\frac{\mathrm{i}}{2EL}\left(1-\frac{(\mu B)^{2}}{2\mathfrak{E}_{\mathrm{B}}^{2}}\right)\right],$ (5.35)
@@
Analogously to Eq. (4.27), Eq. (5.35) is valid for for ultrarelativistic neutrinos. We also keep the terms up to $L^{-2}$ in Eq. (5.35).
@@
where $P_{\mathrm{max}}\approx P_{\nu_{e\mathrm{L}}\to\nu_{\mu\mathrm{L}}}^{(\mathrm{qm})}+P_{\nu_{e\mathrm{L}}\to\nu_{\mu\mathrm{L}}}^{\prime(\mathrm{qft})}+P_{\nu_{e\mathrm{L}}\to\nu_{\mu\mathrm{L}}}^{\prime\prime(\mathrm{qft})}$, and
@@
$P_{\nu_{e\mathrm{L}}\to\nu_{\mu\mathrm{L}}}^{\prime(\mathrm{qft})}$ $=-\frac{(\mu B)^{2}}{E\mathfrak{E}_{\mathrm{B}}}P_{\nu_{e\mathrm{L}}\to\nu_{\mu\mathrm{L}}}^{(\mathrm{qm})},$ (5.39)
$P_{\nu_{e\mathrm{L}}\to\nu_{\mu\mathrm{L}}}^{\prime\prime(\mathrm{qft})}$ $=\frac{1}{4E^{2}L^{2}}\left(1-\frac{(\mu B)^{2}}{2\mathfrak{E}_{\mathrm{B}}^{2}}\right)^{2}P_{\nu_{e\mathrm{L}}\to\nu_{\mu\mathrm{L}}}^{(\mathrm{qm})}.$ (5.40)
@@
The correction to the transition probability $P_{\nu_{e\mathrm{L}}\to\nu_{\mu\mathrm{L}}}^{\prime(\mathrm{qft})}$ in Eq. (5.39) arises from the term $\propto\mu^{2}(\mathbf{B}\cdot\hat{\mathbf{p}})^{2}$ in the propagators in Eqs. (5.32) and (5.33), which is discussed at the end of Sec. 5.3. One can see that $P_{\nu_{e\mathrm{L}}\to\nu_{\mu\mathrm{L}}}^{(\mathrm{qft})}<0$, i.e. the total transition probability is less than the QM prediction. Moreover, one gets from Eqs. (5.38) and (5.39) that $|P_{\nu_{e\mathrm{L}}\to\nu_{\mu\mathrm{L}}}^{\prime(\mathrm{qft})}|/P_{\nu_{e\mathrm{L}}\to\nu_{\mu\mathrm{L}}}^{(\mathrm{qm})}\sim(\mu B)^{2}/\mathfrak{E}_{\mathrm{B}}E\sim\mu B/E\ll 1$ provided that the magnetic contribution is dominant. Therefore, the QFT correction to the propagators, $\propto\mu^{2}(\mathbf{B}\cdot\hat{\mathbf{p}})^{2}$, can be safely neglected since its contribution to the transition probability is negligible. We also mention that the QFT correction to the probability of the spin-flavor precession of Majorana neutrinos, $\nu_{e}\xrightarrow{B}\bar{\nu}_{\mu}$, analogous to that in Eq. (5.39), was obtained in Ref. [18].
The QFT correction to the transition probability in Eq. (5.40), $P_{\nu_{e\mathrm{L}}\to\nu_{\mu\mathrm{L}}}^{\prime\prime(\mathrm{qft})}$ results from keeping the next-to-leading term $\propto L^{-2}$ in Eq. (5.35). This contribution is analogous to the QFT correction derived in Sec. 4.3. One can see that $P_{\nu_{e\mathrm{L}}\to\nu_{\mu\mathrm{L}}}^{\prime\prime(\mathrm{qft})}/P_{\nu_{e\mathrm{L}}\to\nu_{\mu\mathrm{L}}}^{(\mathrm{qm})}=1/16E^{2}L^{2}$ provided that the magnetic contribution is dominant, i.e. $\mathfrak{E}_{\mathrm{B}}\sim\mu B$. Thus, $P_{\nu_{e\mathrm{L}}\to\nu_{\mu\mathrm{L}}}^{\prime\prime(\mathrm{qft})}$ can become important if short macroscopic propagation distances are considered.
@@
One can see in Eq. (5.41) that the leading term reproduces the QM result in Eq. (5.26). A small quantum correction $\sim\mu B/E$ to the amplitude of the survival probability is also present Eq. (5.41). In Eq. (5.41), we omit the QFT correction $\propto L^{-2}$. It can be derived analogously to Eq. (5.40).
6 General issues of the QFT description of neutrino oscillations
Instead of dealing with nondiagonal propagators of neutrino mass eigenstates in external fields, one may consider the ‘propagators’ of neutrino flavor states $S_{\mathrm{F}}$. These ‘propagators’ are $4N_{\nu}\times 4N_{\nu}$ matrices, where $N_{\nu}$ is the number of the neutrino flavors. In vacuum, $S_{\mathrm{F}}$ obeys the modified Dirac equation (see, e.g., Ref. [28]),
$(\mathrm{i}\gamma^{\mu}\partial_{\mu}-M)S_{\mathrm{F}}(x)=\delta^{4}(x),$ (6.1)
where $M$ is the nondiagonal mass matrix; cf. Eq. (2.1). Equation (6.1) is solved in the Fourier representation. The Fourier image of $S_{\mathrm{F}}$ is singular on the mass shell. This singularity is removed by the proper poles bypassing in the standard manner. It is claimed that the poles bypassing gives the meaning of causality to the formal solution of Eq. (6.1).
As we demonstrated in Sec. 3, the probability of neutrino oscillations is based on the element of the $S$-matrix,
$S=\langle\Psi_{\mathrm{out}}|T\exp\bigg[\mathrm{i}\int\mathrm{d}^{4}x\mathcal{L}_{\mathrm{int}}(x)\bigg]|\Psi_{\mathrm{in}}\rangle,$ (6.2)
where $\mathcal{L}_{\mathrm{int}}$ is the interaction Lagrangian and $|\Psi_{\mathrm{in,out}}\rangle$ are the in- and out- Fock states of particles represented as external lines in Fig. 1. In the perturbative approach, the $S$-matrix element corresponds to the Feynman diagram shown in Fig. 1, with neutrinos being represented by the internal line. If one adopts the concept of flavor neutrinos $\nu_{\lambda}$ as fundamental particles, one gets the following two points correlators while decomposing the $S$-matrix element:
$\langle 0_{\mathrm{F}}|\nu_{\lambda}(x)\bar{\nu}_{\lambda^{\prime}}(y)|0_{\mathrm{F}}\rangle=(\tilde{S}_{\mathrm{F}})_{\lambda\lambda^{\prime}}(x-y).$ (6.3)
by using the standard QFT formalism. In Eq. (6.3), $|0_{\mathrm{F}}\rangle$ is the ‘vacuum’ state of flavor neutrinos.
However, it is not obvious that Eq. (6.3), which appears in the calculation of probabilities, coincides with the solution of Eq. (6.1) for a nondiagonal mass matrix $M$: $(\tilde{S}_{\mathrm{F}})_{\lambda\lambda^{\prime}}\overset{?}{=}(S_{\mathrm{F}})_{\lambda\lambda^{\prime}}$. The coincidence of two representations for the causal Green function, or the propagator, can be proven only for mass rather than flavor neutrino eigenstates. That is why, using the solution of Eq. (6.1), and its analogue in the presence of external fields [19, 21], for the calculation of the probabilities is ambiguous.
There is one additional hidden problem with Eq. (6.3). Fock states of an elementary particle correspond to irreducible representations of the Poincaré group [29]. If a particle is massive, each representation is characterized by eigenvalue of the Casimir operator $P^{\mu}P_{\mu}=m_{a}^{2}$. Flavor neutrinos are the superposition mass states, cf. Eq. (2.2). In this sense, all Fock states, including the vacuum one $|0_{\mathrm{F}}\rangle$, of flavor neutrinos are not well defined from the point of view of QFT. Only if all masses of mass states are zero, $m_{a}=0$, the vacua of flavor and mass states are unitary equivalent. Thus, according to the Weinberg definition [30], one can say that a flavor neutrino is not an elementary particle.
On the contrary, the approach proposed in Refs. [13, 14, 15] is based on the neutrino mass eigenstates. Fock states and propagators of such particles are well defined in QFT. In the present work, we generalize the formalism of Refs. [13, 14, 15] to account for external fields. For this purpose, one has to deal with the nontrivial infinite sum of Feynman diagrams in Fig. 2. Nevertheless, in our analysis, we do not go beyond well established methods of QFT.
We use the model of the neutrino emission, propagation and oscillations proposed in Ref. [13]. In this approach, one makes two assumptions: (i) the distance between a source and a detector is constant, and (ii) the incoming and outgoing charged leptons propagate as plane waves, i.e. they have definite momenta. The assumption (i) results in equal energies of charged leptons in Eq. (3.5): $E_{\alpha}=E_{\beta}$. Since masses of these leptons are different, $\mathfrak{m}_{\alpha}\neq\mathfrak{m}_{\beta}$, their momenta momenta are not equal $\mathbf{p}_{\alpha}\neq\mathbf{p}_{\beta}$. To overcome the momentum nonconservation problem, we assume that charged leptons are ultrarelativistic in the present work.
In practice, heavy nuclei in a source and a detector are not at rest. Thus, instead of using the coordinate delta functions in Eq. (3.3), one should integrate over the positions a source and a detector, with the distance between them being kept constant. This procedure was shown in Ref. [31] to be equivalent to the modification of the Feynman propagator. In this case, charged leptons still can be considered as plane waves.
Another possibility is the consideration of the wave packets of charged leptons. That is, $\mathbf{p}_{\alpha,\beta}$ are supposed to be spread around a certain mean momenta $\mathbf{p}_{\alpha,\beta}^{(0)}$. This approach was first proposed in Ref. [14] and, then, developed in numerous works (see, e.g., Ref. [32] for a review). The wave packet formalism allows to solve the energy-momentum conservation problem in neutrino oscillations.
In general, the violation of the above assumptions, either (i) or (ii), leads to the loss of coherence in neutrino oscillations. That is, the oscillatory behavior of the probabilities is suppressed when $L$ exceeds certain value called the coherence length $L_{\mathrm{coh}}$ (see, e.g., Refs. [14, 33]). The computation of $L_{\mathrm{coh}}$ is one of the challenges to a QFT based approach for neutrino oscillations. In a realistic situation, numerous factors, such as the finite lifetime of charged leptons, the energy spread because of measurements etc., contribute to $L_{\mathrm{coh}}$. The recent calculation of $L_{\mathrm{coh}}$ was carried out in Ref. [34]. Note that the value of $L_{\mathrm{coh}}$ can depend on an experimental setup where neutrino oscillations are studied [35].
The majority of calculations of $L_{\mathrm{coh}}$ are made for neutrino oscillations in vacuum. In the vacuum case, even if either the assumption (i) or (ii) is violated, one can still proceed with analytical calculations. When one deals with neutrino oscillations in external fields, the analytical calculation are difficult even in an idealized situation when both assumptions (i) and (ii) are valid (see Appendix B). That is why, the study of decoherence in the QFT approach for neutrino oscillations in external fields, developed in the present work, is rather challenging.
7 Discussion
@@
The technique for finding of the exact propagators is solving the system of the Dyson equations, which are equivalent to summing of infinite series of Feynman diagrams, depicted in Fig. 2. In these diagrams, the nondiagonal interaction with matter is accounted for. This technique was elaborated in Ref. [17] where oscillations of Majorana neutrinos in matter were studied. In case of Dirac neutrinos, the effective matter potential is a singular operator. Since the reciprocal of this operator is involved in the Dyson equations, the straightforward solution of these equations is ambiguous. Nevertheless, we have regularized this potential in Sec. 4.2. After removing the regularization, we have obtained the dressed propagators in background matter in Eqs. (4.17) and (4.20). These propagators have been used in the matrix element to get the transition probability of flavor oscillations in matter in Eq. (4.3). The leading term in Eq. (4.3), $P_{\nu_{e}\to\nu_{\mu}}^{(\text{qm})}$, coincides with the prediction of the QM approach.
@@
Then, we have studied the spin-flavor precession of Dirac neutrinos in a magnetic field in Sec. 5. We have adopted the situation when one has only a transition magnetic moment in the mass basis of two neutrinos. This kind of magnetic moments is disfavored in some models. Nevertheless, such an external magnetic interaction maximally mixes different neutrino mass states. Thus, the QFT formalism for neutrino oscillations based on dressed propagators is highlighted. Moreover, one can consider arbitrary magnetic moments from the phenomenological point of view.
We also mention that the impact of diagonal magnetic moments on the spin-flavor precession of Dirac neutrinos in frames of QFT was previously studied in Ref. [16]. In this situation, it is not required to consider the Dyson Eqs. (4.6) and (4.7) to get nondiagonal dressed propagators. The contribution of the magnetic field can be accounted for by using the exact solution of the Dirac equation for a single neutrino mass eigenstate. The corresponding propagator can be derived analogously to Appendix A, where matter case was studied. On the contrary, if one deals with Dirac neutrinos with a transition magnetic moment, the application of the technique developed in the present work is essential to obtain the dressed propagators in a magnetic field. This fact was the additional motivation to consider this particular type of magnetic moments.
It would be interesting to consider the spin-flavor precession of Dirac neutrinos with an arbitrary matrix of magnetic moments in frames of QFT. However, the Feynman diagrams in Fig. 2, leading to the Dyson Eqs. (4.6) and (4.7), seem to be branching in this situation. The summation of this kind of infinite series is likely to be quite challenging. We plan to study this issue in one of our forthcoming works.
@@
We have revealed a peculiarity of the treatment of the spin-flavor precession of Dirac neutrinos in frames of QFT. This formalism involves both a source, a neutrino propagation, and a detector as inherent parts. Since our model of a detector is based on the standard model physics, where right-handed neutrinos are sterile, we cannot describe the channel like $\nu_{e\mathrm{L}}\to\nu_{\mu\mathrm{R}}$. On the contrary, the process like $\nu_{e\mathrm{L}}\to\nu_{\mu\mathrm{R}}$ can be formally studied in QM description which is provided in Sec. 5.2; cf. Eq. (5.29). In frames of the used model of neutrino oscillations, the spin-flavor precession of active left electron neutrinos is manifest as the deficit of these active particles.
@@
It is the main difference between the Dirac and Majorana cases. In the latter situation, the process like $\nu_{e}\to\bar{\nu}_{\mu}$ is well defined in the QFT formalism. The marker of such a spin-flavor precession is the appearance of an antimuon in a detector.
@@
One can see in the propagators in Eqs. (5.32) and (5.33) that there is a quantum term $\propto\mu^{2}(\mathbf{B}\cdot\hat{\mathbf{p}})^{2}$ which cannot be neglected if we are in frames of QFT. We have derived the matrix element and the transition probability based on these propagators in Sec. 5.4. The leading term in the transition probability in Eq. (5.38) reproduces the QM result in Eq. (5.27). The correction to the QM expression is given in Eq. (5.39). It arises from the quantum term $\propto\mu^{2}(\mathbf{B}\cdot\hat{\mathbf{p}})^{2}$ in the propagators. This correction turns out to be small for ultrarelativistic neutrinos for all reasonable parameters. Thus, we have revealed a posteriori, i.e. after the calculation of a transition probability, that the virtuality of massive neutrinos in propagators is negligible. Note that the correction to the transition probability, analogous to that in Eq. (5.39), was derived in Ref. [18] while applying QFT for the description of the spin-flavor precesion of Majorana neutrinos.
Besides the correction to the transition probabilities of the spin-flavor precession resulting from the QFT contributions to the propagators, we have derived the corrections arising from consideration of the short macroscopic propagation distance. These contributions are $\sim(EL)^{-2}$. The terms in question appear both in the transition probabilities of flavor oscillations in matter and the spin-flavor precession.
When one applies QFT to describe neutrino oscillations, we encounter to some general problems. They include the definition of the Fock states and the propagators of flavor neutrinos, as well as the decoherence in neutrino oscillations. These problems arise even if one studies neutrino oscillations in vacuum. We have discussed these issues in Sec. 6.
@@
We also note that the impact of external fields on neutrino oscillations is an essentially nonperturbative phenomenon which requires the summation of infinite number of Feynman diagrams Fig. 2. Only in this case, one gets the correct form of dressed propagators exactly accounting for an external field, which result in the transition probabilities consistent with the QM predictions.
We also list the main assumptions made in the course of applying of QFT for neutrino oscillations in external fields. The ultrarelativity of neutrinos, the consideration of two mass eigenstates, and the spatial homogeneity of external fields were important in our analysis.
One of the most important assumption made in our work is the ultrarelativity of neutrinos. From the formal point of view, we neglect several terms in the undressed propagators (compare Eq. (A) with Eq. (A.16), as well as Eq. (C) with Eq. (C.2)). Recently, in Ref. [36] we revealed the physical meaning of this kind of procedure. It was demonstrated to be equivalent to disregarding of the antiparticle degrees of freedom. This procedure of the truncation of the undressed propagators allows one to obtain the proper poles in the dressed propagators accounting for the two neutrino states. Note that the truncation of $S_{a}$ can be justified for ultrarelativistic neutrinos only.
However, if one considers the truncated undressed propagators $S_{a}$ in Eqs. (A.16) and Eq. (C.2), as well as omits the terms $\sim m_{a}/E_{a}$ in the numerators of the integrands, $S_{a}$ become singular. While solving the Dyson Eqs. (4.6) and (4.7), one has to deal with the reciprocal of $S_{a}$ in the intermediate computations. That is, one, first, obtains $S_{a}^{-1}$. Then, this quantity is used to get the reciprocal of the dressed propagators, $\Sigma_{ab}^{-1}$. Finally, based on $\Sigma_{ab}^{-1}$, we derive $\Sigma_{ab}$.
Thus, one can guess that the singularity in $S_{a}$ does not affect the final form of $\Sigma_{ab}$ since we invert matrices two times. Nevertheless, certain regularization of $S_{a}$ is required. The same argument is applied to the regularization of the nondiagonal matter potential $G$ in Eq. (4.8). Indeed, one can see in Eq. (4.2) that both $G^{-1}$ and $S_{a}^{-1}$ are infinite when $\alpha=\alpha_{a}=\xi=1$.
One can regularize of the undressed propagators in multiple ways. For example, we could keep the terms linear in $m_{a}$ in the numerators of the integrands in Eqs. (A.16) and Eq. (C.2). However, in this situation, solving Eqs. (4.6) and (4.7) would be more awkward especially in the matter case in Sec. 4.2.
That is why we have chosen an alternative way for the regularization by introducing the parameters $\xi$ and $\alpha_{a}$ in Sec. 4.2, as well as $\varsigma_{a}$ in Sec. 5.3. Note that $\alpha_{a}=(p+g_{a}/2)/E_{a-}$ and $\varsigma_{a}=p/E_{a}$ have the physical meaning. The quantity $\xi$ was introduced formally. All these parameters are supposed to approach to one for ultrarelativistic neutrinos.
We made at least one consistency check for the chosen regularization in Secs. 4.2 and 5.3. We have demonstrated that the diagonal dressed propagators coincide with the undressed ones if one turns off the nondiagonal neutrino interactions, i.e. at $g\to 0$ and $\mu\to 0$ (see the discussion below Eqs. (4.20) and (5.33)).